Imagine you've started a new pizza restaurant. It costs you about $5 to produce a pizza.

Last week you sold 500 pizzas for $10 each.

This week you raised your price and sold 150 pizzas for $16 each.

What price should you be charging?
1
Expert's answer
2020-11-17T12:24:41-0500

Solution:

Derive functions:

"500=(a\times 10) + b"

"150=(a\times 16) + b"

"350=-6a"

"a = \frac{350}{-6}"

"a = -58.3"

Substitute from the function to get b:

"500 = (-58.3\times 10)+b"

"500 = -583+b"

"500+583=b"

"b=1083"

Therefore:

"Q =-58.3P+1083"

Substitute to derive P:

"-58.3P+1083=Q"

"-58.3P=Q-1083"

"P=\frac{-Q}{-58.3}-\frac{1083}{-58.3}"


"P=-0.02Q+18.6"


Find TR:

"TR = (P\times Q)"

"TR=(-0.02Q\times Q) +(18.6\times Q)"

"TR=-0.02Q^{2} +18.6Q"


Find MR:

The MR is the derivative of the total revenue with respect to demand:

Therefore:

"MR =-0.04Q+18.6"


Substitute to derive the Price:

"5=-0.04Q+18.6"

"0.04Q=18.6-5"

"Q=\frac{13.6}{0.04}"


"Q=340"


Substitute for Price:

"P=-0.02Q+18.6"

"P=(-0.02\times 340)+18.6"

"P=-6.8+18.6"

"P=11.8"


The price that should be charged is = 11.8